OpenStudy (anonymous):

. In 2010, the world's population reached 6.91 billion and was increasing at a rate of 1.1% per year. Assume that this growth rate remains constant. (In fact, the growth rate has decreased since 1987.) (a) Write a formula for the world population (in billions) as a function of the number of years since 2010. (b) Estimate the population of the world in the year 2020.

OpenStudy (anonymous):

(a) Okay so the question straight away is a reminder of compound interest: for example imagine a compound interest of 2% on $90. After the first year (x=1) there will be \[90+90\times \frac{ 2 }{ 100 }\] ie (if you factorize)\[90\times1.02\] After the 2nd year (x=2) you will have the same interest rate, but on what you had after the first year: \[(90\times1.02)\times1.02\] which is \[90\times1.02^2\] So we understand that there an underlying formula here\[PV\times(1+r)^n=FV\]:where PV is the present value, r the interest rate, n the number of years for which the interest will be applied (number of periods) and FV our final value. With this in mind, we can easily answer the question \[f(x)=6.91\times10^9\times1.011^x\]where x is the number of years since 2010 (b) Simply input the numbers into the formula, and we get: \[f(20)=8600053623.63\] Now you can't have .63 of a human, so the rational thing to do is to round it. Your final answer is then 8600053624.